jasminegainey77:

An equation was used to predict possible weights of puppies at 6 months of age. The actual weights of the puppies are also listed. Actual weight (ounces) 9 15 16 20 20 26 Predicted weight (ounces) 9 12 15 18 21 24 The sum of the residuals is ______.

justjm:

\(\sum r=\sum (y_i-y_p)\) Hence, you need to take the residuals for each data point, and then add that up. You would take a residual for one data point by subtracting the predicted weight from the actual weight. For instance, for \(x_2\), or month 2, the residual \(r_2=15-12=3\) You can do the rest of the months and then add them up.